Check My Work contacts made during the week. A sample of 55 weekly reports showed a sample mean of 19.5 customer contacts per customer contacts for the sales personnel. The sample standard deviation was 5.9. Provide 90% and 95% confidence intervals for the population mean number of weekly Check My Work
Answer #1

A sample of 55 weekly reports has been chosen from the population. Therefore, the sample size, n = 55.

Now, we have to find 90% and 95% confidence intervals for the population mean. A 90% confidence interval means that this interval will contain the population mean in 90% of the samples of the population, when taken repeatedly.

The formula to calculate the confidence intervals for the population mean is given as: [Ȳ - tα/2,n-1*(s/√n), Ȳ + tα/2,n-1*(s/√n)]

where Ȳ is the sample mean, s is the sample standard deviation and tα/2,n-1 is the value of the T-statistic at the value \alpha.

Why we use a T-statistic here is because we use Student's t distribution when estimating the mean of a normally distributed population in situations where the sample size is small and population standard deviation is unknown. You can use the t-table to find out the values at different values of \alpha.

a. To find 90% confidence interval, \alpha= 1-0.9 = 0.1

tα/2,n-1 = t0.025,54 = 1.67

Ȳ = 19.5

s= 5.9

The confidence interval will be given as : [19.5 - 1.67* (5.9/7.42), 19.5 + 1.67* (5.9/7.42)] = [18.17,20.83]

b. To find 95% confidence interval, \alpha= 1-0.95 = 0.05

tα/2,n-1 = t0.025,54 = 2

Ȳ = 19.5

s= 5.9

The confidence interval will be given as : [19.5 - 2* (5.9/7.42), 19.5 + 2* (5.9/7.42)] = [17.91,21.09]